OpenStudy (anonymous):

3 out of 46 was chosen to do a survey on recommending their attorney. 26 said yes 11 said no 9 said not sure. what would the probability of two out of the chosen 3 would not recommend their attorney and the 3rd one not being sure?

OpenStudy (anonymous):

(11/46)(10/45)(8/44)

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Okay, so first you pick two yes people: \[ \left(\frac{11}{46}\right)\left(\frac{10}{45}\right) \]Then there are 44 left, and you pick a not sure person: \[ \left(\frac{9}{44}\right) \]So to get this order of yes-yes-unsure, you have a probably of: \[ \left(\frac{11}{46}\right)\left(\frac{10}{45}\right)\left(\frac{9}{44}\right) \]Then the question becomes... can we do this another way?

OpenStudy (anonymous):

For example, we could have done unsure-yes-yes. that would give us:\[ \left(\frac{9}{46}\right)\left(\frac{11}{45}\right)\left(\frac{10}{44}\right) \]Notice that no matter the order, the top is still \(11\times 10\times 9\), and the bottom is still \(46\times 45\times 44\).

OpenStudy (anonymous):

So each ordering will still have the same probability. It's just a mater of finding the number of orderings.

OpenStudy (anonymous):

There are three spots. One of them needs to be given 'unsure' and two need to be given yes. This can be counted using combinations. In this case \(\large {3\choose 2}{1\choose 1}\) if you want to do yes then no or \(\large {3\choose 1}{2\choose 2}\). They are both equivalent.

OpenStudy (anonymous):

So the final answer is: \[ {3\choose2}{1\choose1} \left(\frac{11}{46}\right)\left(\frac{10}{45}\right)\left(\frac{9}{44}\right) \] Here is wolfram computation: http://www.wolframalpha.com/input/?i=%7B3%5Cchoose2%7D%7B1%5Cchoose1%7D+%5Cleft%28%5Cfrac%7B11%7D%7B46%7D%5Cright%29%5Cleft%28%5Cfrac%7B10%7D%7B45%7D%5Cright%29%5Cleft%28%5Cfrac%7B9%7D%7B44%7D%5Cright%29

OpenStudy (anonymous):

Oh one minor issue... I said 'yes' when I should have been saying 'no' before.

OpenStudy (anonymous):

so which one is the answer

OpenStudy (anonymous):

The math is still correct though.

OpenStudy (anonymous):

9/44?

OpenStudy (anonymous):

Do you know what the actual answer is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Wolfram calculates it to \(\frac 3{92}\)

OpenStudy (anonymous):

you get the 92 from?

OpenStudy (anonymous):

\[ {3\choose2}{1\choose1} \left(\frac{11}{46}\right)\left(\frac{10}{45}\right)\left(\frac{9}{44}\right) \] Calculating this and simplifying...

OpenStudy (anonymous):

\[ {3\choose 2} = 3, {1\choose 1} = 1 \]\[ \frac{\color{red}{\cancel{11}}\times 10\times \color{blue}{\cancel{9}}}{46\times \color{blue}{\cancel{45}}5\times \color{red}{\cancel{44}}4}=\frac{\color{red}{\cancel{10}}\color{blue}{\cancel{2}}}{46\times \color{red}{\cancel{5}}\times \color{blue}{\cancel{4}}2} = \frac{1}{92} \]So \[ {3\choose 2}{1\choose 1} \frac{11\times 10\times 9}{46\times 45\times 44} = \frac {3}{92} \]

OpenStudy (anonymous):

k I need your help with two more...is that cool?

OpenStudy (anonymous):

dunno, maybe